OpenStudy (anonymous):

Help meeeeeeeeeeeee. Which of the following is the solution to the equation square root of 5^8n = 125^(n + 3) ?

OpenStudy (rsadhvika):

write 125 as 5^3

OpenStudy (anonymous):

okay..then what?

OpenStudy (rsadhvika):

then you will see bases are equal, so exponents also will be equal set them equal and solve n

OpenStudy (anonymous):

can you solve me thrugh it? i get confused on this section real easy..

OpenStudy (rsadhvika):

\(\large \sqrt{5^{8n} } = 125^{(n+3)} \)

OpenStudy (rsadhvika):

\(\large 5^{8n\times \frac{1}{2}} = 5^{3(n+3)} \) \(\large 5^{4n} = 5^{3(n+3)} \)

OpenStudy (rsadhvika):

since bases are equal, you can set the exponents equal

OpenStudy (rsadhvika):

\(4n = 3(n+3) \)

OpenStudy (rsadhvika):

you can solve n now ?

OpenStudy (anonymous):

dude, thats way too long. here is a simpler way 5^8n =125^ n+3 5^8n =5^3(n+3) 8n=3n+9 5n=9 n= 9 / 5